# ref.bib UUITP-22/24

### v1

@@ 3.5.1 $\mathbb{C}^{3}$
Remark 3.5.
In the limit when $\epsilon_{\pm}\to 0$ the leading order behavior of the fourfold quantum volume is dominated by the last line in (3.44)
@@
where $W$ agrees exactly with the equivariant disk potential in [Cassia:2023uvd, eq. (3.28)] after matching conventions (see Remark (2.2)).
@@ 3.5.2 Local $\mathbb{P}^{2}$
In the limit when $\epsilon_{\pm}\to 0$ the leading order behavior of the fourfold quantum volume is obtained by using (see footnote 14)
@@
so that we recover the CY3 quantum volume $\mathcal{F}^{D}(t,\epsilon)$. Note that all dependence on $c$ automatically disappears in the limit.
Taking the limit $\epsilon_{\pm}\to 0$ of the equivariant period associated to the intersection of divisors $D_{\pm}$ we obtain the equivariant disk potential $W(t,c,\epsilon)$ given in [Cassia:2023uvd, Section 5.3] after matching conventions (see Remark (2.2)). In a similar way, we can compute the equivariant periods of the divisors $D_{\pm}$ in $X_{4}$, we leave details to interested readers.
Next, we turn to the extended equivariant Picard–Fuchs equations. The CY4 has both the closed string modulus $t$ and the open string modulus $c$. We can write down operators of the form (3.38). Here the index $a$ takes values $1,2$ corresponding respectively to $(t,c)$, and the extended charge matrix is given in (3.57). Taking $\gamma=(1,0),(0,1)$ and $(1,1)$ gives PF equations for $X_{4}$ involving the open modulus
@@
We next consider the non-equivariant limit of these. Recall the expression (2.47) for the regular part of the monodromy $\Delta_{c}[\mathcal{H}^{D}]_{0}\sim\partial_{c}W_{\mathrm{disk}}\sim\log(-y_{1}(-x))$, with $y_{1}$ given in (2.44). The regular part obeys the following equations
@@
The operators on the left correspond to the naïve non-equivariant limit of the operators in (3.63). The appearance of an inhomogeneous term on the right hand side can be understood by taking the non-equivariant limit carefully. This is generated by the terms of order $O(\epsilon)$ in the operators $\operatorname{\mathcal{L}}_{\gamma}$ acting on the singular terms of order $O(\epsilon^{-1})$ in the equivariant superpotential $W(t,c,\epsilon)$. As in the previous example, we find that the homogeneous extended PF equations are replaced by inhomogeneous equation in the non-equivariant limit. Again this is in line with earlier results from open Gromov–Witten theory, see Remark 3.3.
@@ 4.1 Double cuts from Calabi–Yau fivefolds
Two is the maximum number of symplectic cuts that can be taken simultaneously on a Calabi–Yau threefold while preserving the CY condition. This is because charge vectors of both hyperplanes in the moment polytope of $X_{3}$ must be orthogonal to $(1,1,1)$ and must be linearly independent. Therefore intersections of the hyperplanes always happen along affine lines of slope $(1,1,1)$, and for a generic choice of moduli $c_{\alpha}$, at most two hyperplanes have nontrivial mutual intersection. We define the Calabi–Yau onefold associated to the double cut as follows
$X_{1}:=\left\{\sum_{i=1}^{r+3}Q^{a}_{i}|z_{i}|^{2}=t^{a},\ \sum_{i=1}^{r+3}q^{1}_{i}|z_{i}|^{2}=c_{1},\ \sum_{i=1}^{r+3}q^{2}_{i}|z_{i}|^{2}=c_{2}\right\}\Big{/}U(1)^{r+2}$ (4.1)
Here both $q^{\alpha}$ have entries that add up to zero. Note that this can be viewed either as a standard toric quotient $\mathbb{C}^{r+3}\sslash U(1)^{r+2}$ or as a quotient of the threefold $X_{3}\sslash U(1)^{2}$.
@@
Similar to the fourfold case, $X_{5}$ admits a fibraton over $\mathbb{C}^{2}$
@@
Next we consider fibres of (4.4). If $z_{+}z_{-}=w_{1}$ and $z^{+}z^{-}=w_{2}$ are both nonzero, the fibre is complex isomorphic to $X_{3}$
@@
The fibre degenerates when either of $w_{\alpha}$ vanishes. Over the locus $w_{1}=0,w_{2}\neq 0$ the fibre is complex isomorphic to the symplectic cut defined by $q^{1}$, while over the locus $w_{1}\neq 0,w_{2}=0$ is complex isomorphic to the symplectic cut defined by $q^{2}$. At the origin $w_{1}=w_{2}=0$ the fibre is complex isomorphic to the double cut of $X_{3}$.
We focus on the most degenerate case, i.e. the fibre over the origin. The fibre has now four components corresponding to independent choices of signs in
@@
Each leads to a reduction of (4.2) to a theefold that corresponds to a ‘quarter space’ defined by the double symplectic cut. For example, the reduction of the locus where $z_{-}=z^{-}=0$ corresponds to the intersection of two toric divisors, which we denote as $D_{-}$ and $D^{-}$. The intersection then can be described as
$D_{-}\cap D^{-}=\left\{(x,z_{+},z^{+})\in X_{3}\times\mathbb{C}_{+}\times\mathbb{C}^{+}\;\middle|\;\begin{array}[]{l}\mu_{q^{1}}(x)=c_{1}+|z_{+}|^{2}\,,\\ \mu_{q^{2}}(x)=c_{2}+|z^{+}|^{2}\end{array}\right\}\Big{/}U(1)^{2}\simeq X_{3}^{>,>}\,.$ (4.9)
The remaining three spaces $X_{3}^{>,<},X_{3}^{<,>},X_{3}^{<,<}$ are defined in a similar way.
Recalling that the four spaces $X_{3}^{\lessgtr,\lessgtr}$ are defined by the requirement that $\mu_{q^{1}}(x)\lessgtr c_{1}$ and $\mu_{q^{2}}(x)\lessgtr c_{2}$ in $X_{3}$ (with signs chosen independently), this presentation makes it clear that $X_{3}^{\lessgtr,\lessgtr}$ lift to intersections of toric divisors $D_{\pm}$ and $D^{\pm}$ inside of $X_{5}$.
Of special interest will be the intersection locus $z_{\pm}=z^{\pm}=0$ of the four divisors, which defines a onefold complex isomorphic to $X_{1}$
$D_{+}\cap D_{-}\cap D^{+}\cap D^{-}=\left\{x\in X_{3}\;\middle|\;\mu_{q^{1}}(x)=c_{1}\,,\;\mu_{q^{2}}(x)=c_{2}\right\}/U(1)^{2}\simeq X_{1}\,.$ (4.10)
@@ 4.2 Quantization of the double cut of a CY3
where $x_{\alpha}=\mathrm{e}^{-c_{\alpha}}$ and $z_{a}=\mathrm{e}^{-t^{a}}$ while $A_{\alpha}\cdot\epsilon$ and $t\cdot M\cdot\epsilon$ are certain linear functions of $\epsilon_{i}$ (and $t^{a}$). The function in the denominator is a complex power of $H(x_{1},x_{2},z)$, which is a Laurent polynomial in $x_{\alpha}$ and $z_{a}$. This is related to the Hori–Vafa mirror curve of the Calabi–Yau threefold $X_{3}$ by a simple change of coordinates. More details and a derivation of this formula are given in Appendix C.3.
@@
$\begin{split}\mathcal{H}_{1}^{D}(t,c_{1},\epsilon)&=\int_{\mathbb{R}}\mathop{}\!\mathrm{d}c_{2}\,\mathcal{K}^{D}(t,c_{1},c_{2},\epsilon)\,,\\ \mathcal{H}_{2}^{D}(t,c_{2},\epsilon)&=\int_{\mathbb{R}}\mathop{}\!\mathrm{d}c_{1}\,\mathcal{K}^{D}(t,c_{1},c_{2},\epsilon)\,,\end{split}$ (4.18)
@@ 4.4 Extended Picard–Fuchs equations for double cuts
$\operatorname{\mathcal{L}}_{\gamma}\cdot\mathcal{F}^{D}_{\lessgtr,\lessgtr}(t,c_{1},c_{2},\epsilon)=0\,.$ (4.29)
@@ 4.5.1 $\mathbb{C}^{3}$
The integral can be evaluated in several ways. One of them is to sum over poles selected by an appropriate Jeffrey–Kirwan contour. The choice of contour depends on the phase of the geometry in the moduli space of $c_{1}$, $c_{2}$ (in this case there are no closed string moduli $t^{a}$). Although the contour changes discontinuously with the phase, the quantum volumes obtained by different contours are expected to be related to each other by analytic continuation. See Remark 3.5. Another approach is to switch to integral representations of $\operatorname{\Gamma}$ functions, which leads to $B$-model-like integrals, such as those discussed in Appendix C.
@@
The double cut of the associated CY3 defines a CY1 which we denote as $X_{1}$, whose quantum volume is given by formula (4.12). The computation, whose details are given in Appendix C.3 involves the charge matrix
$\widetilde{Q}=\left[\begin{array}[]{ccc}0&1&-1\\ 1&-1&0\\ \end{array}\right]$ (4.33)
as well as the reduced matrix
$\widetilde{Q}(3)=\left[\begin{array}[]{cc}0&1\\ 1&-1\\ \end{array}\right].$ (4.34)
From these, we obtain the identifications
$A_{1}\cdot\epsilon=\epsilon_{1}+\epsilon_{2}\,,\hskip 30.0ptA_{2}\cdot\epsilon=\epsilon_{1}\hskip 30.0pt\text{and}\hskip 30.0ptH(x_{1},x_{2})=1+x_{1}+x_{1}x_{2}$ (4.35)
through the application of (C.17) and (C.18). Therefore the quantum volume of $X_{1}$ is
@@ 4.5.2 Local $\mathbb{P}^{2}$
This can be evaluated by choosing a JK contour appropriate for the phase under study. The phase structure in this example is more complicated, each phase corresponding to a configuration of hyperplanes for the double symplectic cut, see Figure 8 for some examples. As usual, quantum volumes obtained by different choices of JK contour are expected to be related to each other by analytic continuation, see Remark 3.5.
@@
The double cut of local $\mathbb{P}^{2}$ defines the CY onefold $X_{1}$, with quantum volume given by formula (4.12), corresponding to the charge matrix
$\widetilde{Q}=\left[\begin{array}[]{cccc|c}z_{1}&z_{2}&z_{3}&z_{4}\\ \hline\cr 1&1&1&-3&t\\ 0&0&-1&1&c_{1}\\ 0&1&-1&0&c_{2}\\ \end{array}\right]\,.$ (4.40)
Making use of the reduced matrix (see Appendix (C.3) for definitions)
$\widetilde{Q}(4)=\left[\begin{array}[]{ccc}1&1&1\\ 0&0&-1\\ 0&1&-1\\ \end{array}\right]\,,\qquad\widetilde{Q}(4)^{-1}=\left[\begin{array}[]{ccc}1&2&-1\\ 0&-1&1\\ 0&-1&0\\ \end{array}\right]\,,$ (4.41)
we obtain the following identifications of parameters: $A_{1}\cdot\epsilon=2\epsilon_{1}-\epsilon_{2}-\epsilon_{3}$, $A_{2}\cdot\epsilon=-\epsilon_{1}+\epsilon_{2}$, $t\cdot M\cdot\epsilon=t\epsilon_{1}$ and $H(x_{1},x_{2},z)=1+x_{1}^{-1}+zx_{1}^{2}x_{2}^{-1}+x_{1}^{-1}x_{2}$. The quantum volume of $X_{1}$ is therefore
@@ 5 Summary and conclusions
Our first main result is a general expression for the quantum Lebesgue measure $\mathcal{H}^{D}$ in terms of Lauricella’s hypergeometric function of type $D$, given in (2.24). We then showed that, under certain technical assumptions, the monodromy of its regular part is related to the derivative of the toric brane’s disk potential by formula (2.28). This makes contact with expectations from open string mirror symmetry [Aganagic:2000gs, Aganagic:2001nx] and shows that the correspondence observed in [Cassia:2023uvd] holds in some generality, including in the context of toric geometries not considered there. We expect the correspondence to hold in full generality, and that this may be proven either by studying monodromies of Lauricella’s function or by a careful analysis of the $B$-model-like integral representation for $\mathcal{H}^{D}$ given in (2.22).
Another central point of this work hinges on the description of symplectic cuts in terms of higher-dimensional geometries. We have identified a Calabi–Yau fourfold describing the symplectic cut of a Calabi–Yau threefold, and showed that equivariant periods of the CY4’s divisors and their intersections descend, in a partially non-equivariant limit, to the functions $\mathcal{F}^{D}_{\lessgtr}$ associated to the half-spaces of the symplectic cut, and to the equivariant disk potential $W(t,c,\epsilon)$. An interesting consequence of this correspondence is that the equivariant disk potential must obey extended Picard–Fuchs equations. We have verified this in concrete examples, and found that in the non-equivariant limit the extended PF equations become inhomogeneous.
Finally, we have further generalized our construction by considering double symplectic cuts, which correspond to the maximal allowed number of simultaneous cuts in a Calabi–Yau threefold. We have shown that their geometry is encoded by a Calabi–Yau fivefold, whose equivariant periods correspond to deformations of functions ,$\mathcal{F}^{D}_{\lessgtr,\lessgtr}$ associated with the four subspaces defined by the double cut.
@@
To complete our description of the quantum symplectic cut, it would be desirable to develop an appropriate notion of its quantum cohomology ring. The Calabi–Yau fourfold description of the cut provides a natural approach to this problem, since the CY4 comes with its own Picard–Fuchs equations, and since the various strata of the cut, namely $X_{3}^{\lessgtr}$ and $X_{2}$, arise as toric divisors and intersections thereof. In particular, we have showed that in the limit $\epsilon_{\pm}\to 0$ the CY4 Picard–Fuchs operators descend to extended Picard–Fuchs operators for the CY3 which annihilate $\mathcal{F}^{D}_{\gtrless}$ and the equivariant disk potential $W(t,c,\epsilon)$. One may speculate that these may provide a description of the equivariant quantum cohomology ring for $X_{3}^{<}\cup_{X_{2}}X_{3}^{>}$. It would be interesting to make this expectation more precise. In particular one may ask whether the extended PF equations completely characterize the involved equivariant periods, and what would constitute appropriate initial conditions. Similarly, one may ask whether the extended Picard–Fuchs equations arising from the CY5 can be regarded as a description of the quantum cohomology ring for the double cut of $X_{3}$.
Another question raised in this work is whether double cuts encode more information than just the disk potentials of two toric branes, such as the annulus potential. A natural candidate is the CY1 quantum volume $\mathcal{K}^{D}(t,c_{1},c_{2},\epsilon)$ which depends on the open string moduli of both Lagrangians. It would be interesting to explore whether its analytic structure encodes annuli stratching between branes. A hint that this may be possible comes from the relation to the mirror curve, given in (1.10).

### v2

@@ 3.5.1 $\mathbb{C}^{3}$
In the limit when $\epsilon_{\pm}\to 0$, the leading order behavior of the fourfold quantum volume is dominated by the last line in (3.44)
@@
where $W$ agrees exactly with the equivariant disk potential in [Cassia:2023uvd, eq. (3.28)] after matching conventions (see Remark 2.1).
@@ 3.5.2 Local $\mathbb{P}^{2}$
In the limit when $\epsilon_{\pm}\to 0$ the leading order behavior of the fourfold quantum volume is obtained by using (see footnote 16)
@@
so that, at leading order in ϵ±subscriptitalic-$\epsilon_{\pm}$, we recover the CY3 quantum volume $\mathcal{F}^{D}(t,\epsilon)$ as in Section 2.3.2. Note that all dependence on $c$ automatically disappears in the limit.
Taking the limit $\epsilon_{\pm}\to 0$ of the equivariant period $\Pi(D_{+}\cap D_{-})$ associated to the intersection of divisors $D_{\pm}$ according to the general formula (3.19), we obtain the equivariant disk potential $W(t,c,\epsilon)$ given in [Cassia:2023uvd, Section 5.3] after matching conventions (see Remark 2.1). We leave the details of the computation to interested readers.
Next, we turn to the extended equivariant Picard–Fuchs equations. The CY4 has both the closed string modulus $t$ and the open string modulus $c$. We can write down operators of the form (3.38). Here the index $a$ takes values $1,2$ corresponding, respectively, to $(t,c)$, and the extended charge matrix is given in (3.57). Taking $\gamma=(1,0),(0,1)$ and $(1,1)$ gives PF equations for $X_{4}$ involving the open modulus
@@
We next consider the non-equivariant limit of these PF equations. Recall expression (2.51) for the regular part of the monodromy $\Delta_{c}[\mathcal{H}^{D}]_{0}\propto\partial_{c}W_{\mathrm{disk}}\propto\log(-y_{1}(-x))$, with $y_{1}$ given in (2.48). The regular part obeys the following equations
@@
The operators on the left correspond to the naïve non-equivariant limit of the operators in (3.63), while the inhomogeneous term on the r.h.s. appears in the same way it did for $\mathbb{C}^{3}$ in the last section. This is generated by the terms of order $O(\epsilon)$ in the operators $\operatorname{\mathcal{L}}_{\gamma}$ acting on the singular terms of order $O(\epsilon^{-1})$ in the equivariant disk potential $W(t,c,\epsilon)$. As in the previous example, we find that the homogeneous extended PF equations are replaced by inhomogeneous equation in the non-equivariant limit. Again this is in line with earlier results from open Gromov–Witten theory, see Remark 3.2.
@@ 4.1 Double cuts from Calabi–Yau fivefolds
Similar to the fourfold case, $X_{5}$ admits a fibration over $\mathbb{C}^{2}$
@@
Next we consider fibers of (4.3). If $z_{+}z_{-}=w_{1}$ and $z^{+}z^{-}=w_{2}$ are both nonzero, the fiber is complex isomorphic to $X_{3}$
@@
The fiber degenerates when either of $w_{\alpha}$ vanishes. Over the locus $w_{1}=0,w_{2}\neq 0$ the fiber is complex isomorphic to the symplectic cut defined by $q^{1}$, while over the locus $w_{1}\neq 0,w_{2}=0$ is complex isomorphic to the symplectic cut defined by $q^{2}$. At the origin $w_{1}=w_{2}=0$ the fiber is complex isomorphic to the double cut of $X_{3}$.
We focus on the most degenerate case, i.e. the fiber over the origin. The fiber has now four components corresponding to independent choices of signs in
@@
Each leads to a reduction of (4.1) to a threefold that corresponds to a ‘quarter space’ defined by the double symplectic cut. For example, the reduction of the locus where $z_{-}=z^{-}=0$ corresponds to the intersection of two toric divisors, which we denote as $D_{-}$ and $D^{-}$. The intersection can be described as
$D_{-}\cap D^{-}=\left\{(x,z_{+},z^{+})\in X_{3}\times\mathbb{C}_{+}\times\mathbb{C}^{+}\;\middle|\;\begin{array}[]{l}\mu_{q^{1}}(x)=c_{1}+|z_{+}|^{2}\,,\\ \mu_{q^{2}}(x)=c_{2}+|z^{+}|^{2}\end{array}\right\}\Big{/}U(1)^{2}\,,$ (4.8)
which we will denote as $X_{3}^{>,>}$ to indicate that it is obtained as a reduction of the locus where $\mu_{q^{1}}(x)>c_{1}$ and $\mu_{q^{2}}(x)>c_{2}$. The remaining three spaces $X_{3}^{>,<},X_{3}^{<,>},X_{3}^{<,<}$ are defined in a similar way by the requirement that $\mu_{q^{1}}(x)\lessgtr c_{1}$ and $\mu_{q^{2}}(x)\lessgtr c_{2}$ in $X_{3}$ (with signs chosen according to the labels). This presentation makes it clear that all of the $X_{3}^{\lessgtr,\lessgtr}$ are obtained as intersections of toric divisors $D_{\pm}$ and $D^{\pm}$ inside of $X_{5}$.
Of special interest will be the intersection locus $z_{\pm}=z^{\pm}=0$ of all four divisors, which defines a complex variety of dimension one
$D_{+}\cap D_{-}\cap D^{+}\cap D^{-}=\left\{x\in X_{3}\;\middle|\;\mu_{q^{1}}(x)=c_{1}\,,\;\mu_{q^{2}}(x)=c_{2}\right\}/U(1)^{2}\,.$ (4.9)
Note that, by a standard result for symplectic quotients known as “reduction in stages[MR2091310], this variety can be viewed either as a quotient of the threefold $X_{3}\sslash U(1)^{2}$ or equivalently as the quotient $\mathbb{C}^{r+3}\sslash U(1)^{r+2}$. We will then define the onefold associated to the double cut as follows
$X_{1}:=\left\{\sum_{i=1}^{r+3}Q^{a}_{i}|z_{i}|^{2}=t^{a},\ \sum_{i=1}^{r+3}q^{1}_{i}|z_{i}|^{2}=c_{1},\ \sum_{i=1}^{r+3}q^{2}_{i}|z_{i}|^{2}=c_{2}\right\}\Big{/}U(1)^{r+2}\,.$ (4.10)
Here both $q^{\alpha}$ have entries that add up to zero because of our assumptions on the type of cuts, which implies that the canonical bundle of $X_{1}$ is trivial, so that it is a Calabi–Yau onefold.
@@ 4.2 Quantization of the double cut of a CY3
where $x_{\alpha}=\mathrm{e}^{-c_{\alpha}}$ and $z_{a}=\mathrm{e}^{-t^{a}}$ while $A_{\alpha}\cdot\epsilon$ and $t\cdot M\cdot\epsilon$ are certain linear functions of $\epsilon_{i}$ (and $t^{a}$). The function in the denominator is a complex power of $H(x_{1},x_{2},z)$, which is a Laurent polynomial in $x_{\alpha}$ and $z_{a}$. It is related to the Hori–Vafa mirror curve of the Calabi–Yau threefold $X_{3}$ by a simple change of coordinates. More details and a derivation of this formula are given in Appendix B.3.
@@
$\begin{split}\mathcal{H}_{1}^{D}(t,c_{1},\epsilon)&=\int_{\mathbb{R}}\mathop{}\!\mathrm{d}c_{2}\,\mathcal{K}^{D}(t,c_{1},c_{2},\epsilon)\,,\\ \mathcal{H}_{2}^{D}(t,c_{2},\epsilon)&=\int_{\mathbb{R}}\mathop{}\!\mathrm{d}c_{1}\,\mathcal{K}^{D}(t,c_{1},c_{2},\epsilon)\,.\end{split}$ (4.18)
@@ 4.4 Extended Picard–Fuchs equations for double cuts
$\left.\operatorname{\mathcal{L}}_{\gamma}\right|_{\epsilon_{\pm}=0}\cdot\mathcal{F}^{D}_{\lessgtr,\lessgtr}(t,c_{1},c_{2},\epsilon)=0\,.$ (4.29)
@@ 4.5.1 $\mathbb{C}^{3}$
The integral can be evaluated in several ways. One of them is to sum over poles selected by an appropriate Jeffrey–Kirwan contour. The choice of contour depends on the phase of the geometry in the moduli space of $c_{1}$, $c_{2}$ (in this case there are no closed string moduli $t^{a}$). Although the contour changes discontinuously with the phase, the quantum volumes obtained by different contours are expected to be related to each other by analytic continuation. See Remark 2.2.
@@
Another approach is to switch to integral representations of $\operatorname{\Gamma}$ functions, which leads to $B$-model-like integrals, such as those discussed in Appendix B.
The double cut of the associated CY3 defines a CY1 which we denote as $X_{1}$, whose quantum volume is given by formula (4.12), which in this case reads
@@ 4.5.2 Local $\mathbb{P}^{2}$
The phase structure in this example is more complicated, each phase corresponding to a configuration of hyperplanes for the double symplectic cut, leading to different choices of contours for the integral. See Figure 8 for some examples of possible phases.
@@
This double cut of local $\mathbb{P}^{2}$ defines a CY onefold $X_{1}$ with quantum volume given by
@@ 5 Summary and conclusions
Our first main result is a general expression for the quantum Lebesgue measure $\mathcal{H}^{D}$ in terms of Lauricella’s hypergeometric function of type $D$, given in (2.30). We then showed that, under certain technical assumptions, the monodromy of its regular part is related to the derivative of the toric brane’s disk potential by formula (2.32). This makes contact with expectations from open string mirror symmetry [Aganagic:2000gs, Aganagic:2001nx] and shows that the correspondence observed in [Cassia:2023uvd] holds in larger generality, including in the context of toric geometries not considered there. We expect the correspondence to hold for all (CY) symplectic cuts of both smooth and orbifold toric Calabi–Yau threefolds, and that this may be proven either by studying monodromies of Lauricella’s function or by a careful analysis of the $B$-model-like integral representation for $\mathcal{H}^{D}$ given in (2.28).
Another central point of this work hinges on the description of symplectic cuts in terms of higher-dimensional geometries. Based on Braverman’s observation that the symplectic cut of a Calabi–Yau threefold can be described via an associated Calabi–Yau fourfold, we have showed that equivariant periods of the CY4’s divisors and their intersections descend, in a partially non-equivariant limit, to the functions $\mathcal{F}^{D}_{\lessgtr}$ associated to the half-spaces of the symplectic cut, and to the equivariant disk potential $W(t,c,\epsilon)$. An interesting consequence of this correspondence is that the equivariant disk potential must obey extended Picard–Fuchs equations. We have verified this in concrete examples, and found that in the non-equivariant limit the extended PF equations become inhomogeneous.
Finally, we have further generalized our construction by considering double symplectic cuts, which correspond to the maximal allowed number of simultaneous CY cuts of a Calabi–Yau threefold. We have shown that their geometry is encoded by a Calabi–Yau fivefold, whose equivariant periods correspond to deformations of functions ,$\mathcal{F}^{D}_{\lessgtr,\lessgtr}$ associated with the four subspaces defined by the double cut.
@@
To complete our description of the quantum symplectic cut, it would be desirable to develop an appropriate notion of its quantum cohomology ring. The Calabi–Yau fourfold description of the cut provides a natural approach to this problem, since the CY4 comes with its own Picard–Fuchs equations, and since the various strata of the cut arise as toric divisors as in the case of $X_{3}^{\lessgtr}$, or as intersections of divisors as in the case of $X_{2}$. In particular, we have showed that in the limit $\epsilon_{\pm}\to 0$ the CY4 Picard–Fuchs operators descend to extended Picard–Fuchs operators for the CY3 which annihilate $\mathcal{F}^{D}_{\gtrless}$ and the equivariant disk potential $W(t,c,\epsilon)$. One may speculate that these may provide a description of the equivariant quantum cohomology ring for $X_{3}^{<}\cup_{X_{2}}X_{3}^{>}$. It would be interesting to make this expectation more precise. In particular one may ask whether the extended PF equations completely characterize the involved equivariant periods, and what would constitute appropriate initial conditions. Similarly, one may ask whether the extended Picard–Fuchs equations arising from the CY5 can be regarded as a description of the quantum cohomology ring for the double cut of $X_{3}$.
A related question is how to interpret the symplectic cut procedure from the perspective of local mirror symmetry. In particular, it would be interesting to clarify the role of the CY fourfold $X_{4}$ on the mirror-dual side. We expect that manipulations of equivariant $B$-model integrals, such as those discussed in Appendix B, may provide insight into this direction.
Another question raised in this work is whether double cuts encode more than just the disk potentials of two toric branes—specifically, whether they also capture the annulus potential. A natural candidate for this additional information is the CY1 quantum volume $\mathcal{K}^{D}(t,c_{1},c_{2},\epsilon)$, which depends on the open string moduli of both Lagrangians. It would be interesting to investigate whether its analytic structure encodes annuli stretching between branes. A hint that this might be the case comes from its relation to the mirror curve, as given in (1.10). A direct approach might be to study the GLSM on a circle with suitable boundary conditions at the two spatial boundaries, suggesting that candidates for the annulus potential might be the Witten index, or certain generalizations thereof [Cecotti:1992qh]. [^21] We thank the anonymous referee for suggesting a possible role of the Witten index in relation to annulus potentials. It would be very interesting to explore this question further, but we leave it to future work.